HUBBLE SPACE TELESCOPE DAILY REPORT #2237 PERIOD COVERED: 0000Z (UTC) 10/26/98 - 0000Z (UTC) 10/27/98 Daily Status Report as of 300/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 7590 (Cosmological Parameters Omega and Lambda From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova 5H1.3ZA and 23H1.3ZA in order to study its redshift. The HST was used to make direct measurements of the cosmological parameters Omega, Lambda, and curvature using Type 1A supernovae as a calibrated "standard candle". The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Eight Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.5 Completed Five Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation II) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Six Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed NIC/1/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the young galaxy Q0102-0214 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAs). The DLA line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/MA2 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.9 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star LHS445. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.10 Completed STIS/CCD/MA1/MA2 7434 (The Supernova Intensive Study) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to continue the long-term project on supernovae by observing SN1998S. HST can make unique contributions to understanding supernovae, as has been demonstrated through this program in past years, and now STIS and NICMOS offer new opportunities for deepening this investigation. The aim of the proposal is to explore the UV emission from supernovae, exploit the spatial resolution of HST and press the late-time observations of supernovae into uncharted territory. The observations were completed as planned, and no anomalies were reported. 1.11 Completed Three Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.14 Completed STIS/CCD 8080 (STIS Spectroscopic CCD Flatfielding -- Revised) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat fielding calibration observations in the spectroscopic mode. The observations were completed as planned, and no anomalies were reported. 1.15 Completed STIS/CCD/MA1/MA2 7389 (Chromospheres and transition layers in Hyades and Pleiades F stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the F class star HD29225 which is located in the Hyades star cluster. This proposal seeks to obtain additional empirical constraints on the heating mechanisms for stellar chromospheres, transition layers and coronae. A wavelenth calibration obseration was also taken. The observations were completed as planned, and no porblems were reported. 1.16 Completed FGS 7677 (Monitor the Long Term Stability of FGS-3) The goal of these measurements is to monitor the internal positional stability of FGS3 by measuring many stars in the M35 check field repeatedly over a one year period. The results of these measurements are the stability characteristics of FGS3 which allow an estimate of differential corrections to OFAD and plate scale. The measurements were completed as scheduled. with no reported anomalies. 1.17 Completed WF/PC-2 6659 (A Search For Superplanets Around Weak T- Tauri Stars) The WF/PC-2 was used to make observations in the 1 micron band of the T-Tauri class star HBC426 in order to search for superplanets. Weak T-Tauri class stars are thought to have dissipated their circumstellar disks recently, possibly by forming planetary systems. The observations were completed as planned, and no anomalies were reported. 1.18 Completed WF/PC-2 7376 (Understanding Halo Microlensing and Variable Stars in M31: 1} Snapshot Survey of Candidates) The WF/PC-2 was used to make observations of the galaxy M-31-Lens system #4 to help understand its halo microlensing and Variable Stars. Six microlensing candidates affecting stars in M31 and presumably due to sub-solar-mass MACHOs are being studied. Simultaneously we can explored the stellar population in the relatively unexplored disk/bulge transition via its variable stars, luminosity function and color-magnitude diagram. The observations were executed as scheduled, and no anomalies were reported. 1.19 Completed NIC/1 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera One) was used to make narrow band observations of the BL-Lac (BAL) Q2154-2005. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 7 The acquisition at 300/041259Z failed when the scan step limit was exceeded, indicating a likely walkdown failure. HSTAR 6740 was written. One proposal, to be detailed in tomorrow's report, was affected: NICMOS 7894. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: The DF-224 vehicle clock rollover (to the 32nd most significant bit) occurred yesterday at approximately 299/204215Z. The event was nearly nominal, with only the Thermal Management System software being affected. That software is expected to be brought back on-line today. The SSR EDAC error counter was cleared once per ROP SR-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH